OpenStudy (anonymous):

Solve B = –4(c – 2d) for d Please explain!

OpenStudy (anonymous):

to solve for d you want d alone in one side so -b/4=c-2d c-b/4=2d d=(c-b/4)/2 or d= (4c-b)/8

OpenStudy (anonymous):

My options are in fraction form Would is be d=b-4c/8 ?

OpenStudy (anonymous):

it should be like that d= -(b-4c)/8 ,if i didn't do anything wrong

OpenStudy (anonymous):

Okay thank you!

OpenStudy (anonymous):

u r welcome